# Evaluating the Effectiveness of LLMs in Introductory Computer Science Education: A Semester-Long Field Study

### v2

@@ Abstract.
copyright: acmcopyrightconference: In Proceedings of the Tenth ACM Conference on Learning @ Scale; July 18-July 20, 2024; Atlanta, Georgia, GA, USAjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: In Proceedings of the Tenth ACM Conference on Learning @ Scale; July 18–20, 2024; Atlanta, Georgia, GA, USAbooktitle: L@S’24:In Proceedings of the Tenth ACM Conference on Learning @ Scale, July 18–20, 2024, Atlanta, Georgia, GA, USAprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
The unique challenges posed by LLMs stem from the difficulty in detecting the use of AI tools (Wu et al., 2023b). Traditional approaches, such as plagiarism detection software, fall short in determining the originality of student submissions (Meyer et al., 2023). Given the challenges in identifying LLMs usage and recognizing the potential advantages of these technologies, we consider integrating LLMs into computer science education inevitable. As students have already started using such tools, the impact of LLMs on computer science education remains unknown. Indeed, a growing body of research has begun to explore the application of LLMs within educational settings, primarily focusing on assessing the capabilities of current models with existing datasets or previous assignments from students (Hicke et al., 2023; Mehta et al., 2023). However, there is still a research gap in understanding how students interact with LLM-powered tools in introductory programming classes, particularly regarding their engagement in genuine learning settings over extended periods. Furthermore, while previous studies have shown individual differences in intelligent tutoring systems (Kulik and Fletcher, 2016), research into how these differences apply to LLM tools is lacking. Investigating these variations is important for tailoring educational strategies to diverse student needs. In short, understanding these nuanced attitudes of and interactions with LLM-powered tools in CS education over extended periods is crucial for identifying the evolving challenges and opportunities LLMs introduce.
@@ 2.2. Large Language Models in CS Education
Furthermore, prior research has demonstrated that individual differences, such as gender, race, and prior experiences with technologies, significantly influence the effectiveness of traditional intelligent tutoring systems (Kulik and Fletcher, 2016). However, work that examines how these individual differences affect interactions with and perceptions of LLM-powered tools in educational settings is sparse. Given the increasing integration of LLMs in programming courses and beyond, understanding the role of demographic and individual variability is crucial for developing inclusive and effective educational tools that suit diverse students’ needs.
@@ 3.2. Participants
Race White 26
@@ 4.1.2. Effect of CodeTutor Accessibility on Academic Performance
By constructing the CodeTutor-Allowed and CodeTutor-Not-Allowed, we determine the correlation between CodeTutor’s accessibility and student academic performance. Using the ANOVA technique on the data from the experiment group, Figure 3 reveals that the mean score for the CodeTutor-Allowed category stands at 102.29, in contrast to the CodeTutor-Not-Allowed components, which has a mean score of 93.40. The statistical analysis results show a significant difference between the two groups ($t=$ 2.31, $p=$ 0.03), suggesting that the allowance of CodeTutor correlates with higher student scores.
@@ 4.1.3. Correlation Between Student Demographics and Final Scores in the Experimental Group
Table 2. Multiple regression models explaining respondents’ final scores in experiment group. (Significance level: $p<0.1$, * $p<0.05$, ** $p<0.01$, *** $p<0.001$).
@@
Subsequently, we evaluated demographic factors to determine whether specific student groups, particularly those with prior tech experience, experienced greater benefits from CodeTutor. subsubsection 4.1.3 shows the results of multiple regression models, examining how students’ final scores in the experiment group are associated with their LLM history, major, gender, and race. Students who have never used any LLM-powered tools performed a significant increase ($\beta=$ 18.877, $p=$ 0.032) in final score than the students who used it before.
Moreover, differences in final scores among various majors within the experiment group were statistically significant, indicating that majors play a substantial role in final scores in the experiment group. Students majoring in data science ($\beta=$ 14.532, $p=$ 0.073), mathematics ($\beta=$ 17.692, $p=$ 0.057), and biology ($\beta=$ 16.257, $p=$ 0.057) exhibited a significant positive correlation with final scores compared to those majoring in computer science, suggesting that these majors achieved higher final scores. In terms of gender, no significant effects were observed, indicating no difference between genders in final scores. Additionally, no significant differences were noted across the races in final scores.
@@ 4.2.1. Descriptive Analysis
In terms of students’ attitudes towards CodeTutor (see Figure 1 [Uncaptioned image] for the specific questions), we found that a small portion of students (8%) strongly disagreed or disagreed that CodeTutor accurately understood what students intended to ask, while most (67%) agreed or strongly agreed. In addition, 35% strongly disagreed or disagreed that CodeTutor helped them think critically, while 19% agreed or strongly agreed. Furthermore, 13% students disagreed that CodeTutor improved their understanding of programming syntax, with a larger proportion of individuals agreeing (33%) or strongly agreeing (25%). Nearly half of the students (42%) agreed or strongly agreed that CodeTutor helped students build their own understandings, while very few (17%) strongly disagreed or disagreed. Finally, regarding the potential of CodeTutor to substitute for a human teaching assistant, 20% of the students strongly disagreed or disagreed with this notion, while 42% of them agreed or strongly agreed. Figure 4 shows the distribution of students’ responses across these five questions.

### v3

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the Eleventh ACM Conference on Learning @ Scale; July 18–20, 2024; Atlanta, GA, USAbooktitle: Proceedings of the Eleventh ACM Conference on Learning @ Scale (L@S ’24), July 18–20, 2024, Atlanta, GA, USAdoi: 10.1145/3657604.3662036isbn: 979-8-4007-0633-2/24/07ccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
The unique challenges posed by LLMs stem from the difficulty in detecting the use of AI tools (Wu et al., 2023b; Zhou et al., 2023). Traditional approaches, such as plagiarism detection software, fall short in determining the originality of student submissions (Meyer et al., 2023). Given the challenges in identifying LLMs usage and recognizing the potential advantages of these technologies, we consider integrating LLMs into computer science education inevitable. As students have already started using such tools, the impact of LLMs on computer science education remains unknown. Indeed, a growing body of research has begun to explore the application of LLMs within educational settings, primarily focusing on assessing the capabilities of current models with existing datasets or previous assignments from students (Hicke et al., 2023; Mehta et al., 2023). However, there is still a research gap in understanding how students interact with LLM-powered tools in introductory programming classes, particularly regarding their engagement in genuine learning settings over extended periods. Furthermore, while previous studies have shown individual differences in intelligent tutoring systems (Kulik and Fletcher, 2016), research into how these differences apply to LLM tools is lacking. Investigating these variations is important for tailoring educational strategies to diverse student needs. In short, understanding these nuanced attitudes of and interactions with LLM-powered tools in CS education over extended periods is crucial for identifying the evolving challenges and opportunities LLMs introduce.
@@ 2.2. Large Language Models in CS Education
Furthermore, prior research has demonstrated that individual differences, such as gender, race, and prior experiences with technologies, significantly influence the effectiveness of intelligent tutoring systems (Kulik and Fletcher, 2016). However, work that examines how individual differences affect interactions with and perceptions of LLM-powered tools in educational settings is sparse, even though understanding the role of demographic and individual variability is crucial (Zhang et al., 2024). This is particularly important for developing inclusive and effective educational tools that suit the diverse needs of students.
@@ 3.2. Participants
Race African American or Black 1
@@ 4.1.2. Effect of CodeTutor Accessibility on Academic Performance
By constructing the CodeTutor-Allowed and CodeTutor-Not-Allowed, we determine the correlation between CodeTutor’s accessibility and student academic performance. Using the ANOVA technique on the data from the experimental group, Figure 3 reveals that the mean score for the CodeTutor-Allowed category stands at 102.29, in contrast to the CodeTutor-Not-Allowed components, which has a mean score of 93.40. The statistical analysis results show a significant difference between the two groups ($t=$ 2.31, $p=$ 0.03), suggesting that the allowance of CodeTutor correlates with higher student scores.
@@ 4.1.3. Correlation Between Student Demographics and Final Scores in the Experimental Group
Table 2. Multiple regression models explaining respondents’ final scores in experimental group. (Significance level: $p<0.1$, * $p<0.05$, ** $p<0.01$, *** $p<0.001$).
@@
Subsequently, we evaluated demographic factors to determine whether specific student groups, particularly those with prior tech experience, experienced greater benefits from CodeTutor. subsubsection 4.1.3 shows the results of multiple regression models, examining how students’ final scores in the experimental group are associated with their LLM history, major, gender, and race. Students who have never used any LLM-powered tools performed a significant increase ($\beta=$ 18.877, $p=$ 0.032) in final score than the students who used it before.
Moreover, differences in final scores among various majors within the experimental group were statistically significant, indicating that majors play a substantial role in final scores in the experimental group. Students majoring in data science ($\beta=$ 14.532, $p=$ 0.073), mathematics ($\beta=$ 17.692, $p=$ 0.057), and biology ($\beta=$ 16.257, $p=$ 0.057) exhibited a significant positive correlation with final scores compared to those majoring in computer science, suggesting that these majors achieved higher final scores. In terms of gender, no significant effects were observed, indicating no difference between genders in final scores. Additionally, no significant differences were noted across the races in final scores.
@@ 4.2.1. Descriptive Analysis
In terms of students’ attitudes towards CodeTutor (see Figure 1 [Uncaptioned image] for the specific questions), we found that a small portion of students (8%) strongly disagreed or disagreed that CodeTutor accurately understood what students intended to ask, while most (67%) agreed or strongly agreed. In addition, 35% strongly disagreed or disagreed that CodeTutor helped them think critically, while 19% agreed or strongly agreed. Furthermore, 13% students disagreed that CodeTutor improved their understanding of programming syntax, with a larger proportion of individuals agreeing (33%) or strongly agreeing (25%). Nearly half of the students (42%) agreed or strongly agreed that CodeTutor helped students build their own understandings, while very few (17%) strongly disagreed or disagreed. Finally, regarding the potential of CodeTutor to substitute for a human teaching assistant [^1] In our analysis, response values to the TA Replacement question were reversed, so a higher score indicates a stronger preference for our tool over human teaching assistants. This reversal is consistently applied across all subsequent analyses., 20% of the students strongly disagreed or disagreed with this notion, while 42% of them agreed or strongly agreed. Figure 4 shows the distribution of students’ responses across these five questions.